175Q16 Because the number of moles of gas are increasing from 6 to 7 in the...

175Q16

Because the number of moles of gas are increasing from 6 to 7 in the reaction shown below, at constant pressure ΔE is predicted to be slightly ________ negative than ΔH.

C3H8(g) + 5 O2(g) → 3 CO2(g) + 4 H2O(g)
Answer #1

C3H8(g) + 5 O2(g) → 3 CO2(g) + 4 H2O(g)

Δn = total no of moles of products - total no of moles of reactants in gases

     = (3 + 4) - (1 + 5) = 1

ΔH    = ΔE + ΔnRT

ΔH    = ΔE + 1RT

ΔE   = ΔH -RT

at constant pressure ΔE is predicted to be slightly RT negative than ΔH.